# A Physics-Constrained Neural Differential Equation Framework for Data-Driven Snowpack Simulation

### v1

@@ 2.1 Overview
Our model choices leverage ML for seasonal snow simulation in climate models, prioritizing generalizability and computational efficiency. Contemporary paradigms in hydrology research focus on parameterizing SWE from $z$ and other data to constrain its value over global grids. However, within global climate models, SWE evolution is already well-constrained by explicitly implemented physical laws enforcing mass conservation and relatively well-understood fluxes such as sublimation, precipitation, and melt. The subsequent conversion of these SWE values to variables like $z$ is typically left to longstanding parameterizations. Such models are then calibrated to non-SWE data or even gridded SWE estimates. The resulting simulations are sometimes employed in refining gridded SWE for calibrating other models, creating circular estimation and biases.
@@ 2.2 Model Structure
Figure 1: Structure of the model’s predictive component. Blue lines indicate a trainable linear transformation of the input (of $k$ scalar variables), including a bias. Colors indicate the activation function used upon collection at the node, as noted in the legend. The hyperparameter $n$ sets the width of the internal mixing layer.
@@ 2.3 Threshold Constraints for Snowpack Prediction
Constraints for $dz/dt$ should keep the depth tendency within physical limits to enable generalizability and stability when $M$ is integrated over time. We chose the following constraints:
@@
These constraints can be expressed as threshold functions, the lower as $f_{-}=-z/\Delta t$ and the upper as $f_{+}={\rm{ReLU}}(p)\times\mathbf{1}_{P_{\mathrm{snow}}>0}$, where $p$ is the output of the predictive component. For these choices, $f_{-}$ is nonpositive and $f_{+}$ is nonnegative, which simplifies the constraint layer structure (see Appendix A), resulting in a final structure for $M$ as depicted in Fig. 2.
@@
The first constraint includes the time step $\Delta t$, but this does not explicitly affect the time dependency or resolution of the parameterization. The predictive component contains no time nor time-step dependence. Adjusting $\Delta t$ scales the constraint appropriately without altering the predictive component’s output, enabling the model’s use in adaptive time-step schemes. This means the model in principle only requires training with data at one temporal resolution, though we would anticipate improved time-step independence if training data incorporated varied time intervals, or $dz/dt$ values in the minimum range anticipated during usage.
@@ 2.4 Data
We required training data with simultaneous snow and meteorological measurements, preferring collocated primary ground observations over reanalysis data due to known discrepancies (Meyer et al., 2023). One primary source is the US Snow Telemetry Network (SNOTEL) by the National Resources Conservation Service (NRCS). Data from 44 total SNOTEL sites in the contiguous United States (CONUS) have simultaneous availability of the necessary inputs, of which the entire reporting history until February 1st, 2024 was collected. These sites span diverse climates (see Fig 3), enhancing generalizability. For testing, 7 total Alaskan SNOTEL sites had the necessary inputs and were similarly collected, plus standard evaluation data from Kühtai, Austria (Krajči et al., 2017), Col de Porte, France (Lejeune et al., 2019), and the Reynolds Mountain East catchment (Reba et al., 2011) to assess the $M$’s ability to generalize to out-of-sample data. We additionally used data from Sodankylä, Finland (Essery et al., 2016), the upper Rofental (Warscher et al., 2024), and Yala Basecamp (Stigter et al., 2021; Shea et al., 2015) in the Himalayas to test performance across uncalibrated elevations and climate types.
@@ 2.6 Testing
The integer $K$ enabled traversal of data “gaps” (from missing or cleaned data) larger than $\Delta t=$ 1 day, with $K=1$ for sequential transitions. Predictions over gaps with $K>1$ producing negative $z$ values were set to zero, and timeseries were “reset” to observed values $\hat{z}_{i+K}=z_{i+K}$ if $K>K_{\mathrm{max}}=5$ days. This reduced resets that could otherwise skew performance, with $K_{\mathrm{max}}=5$ ensuring resets in less than 2.3% of cases (median frequency 0.19%), with many over whole years or no-snow periods (not impacting snow simulation). Evaluation metrics included root mean square error (RMSE), mean absolute error (MAE), bias (B), and median percent error (MPE, on $dz/dt$ for regression and $z$ for timeseries). For timeseries, we included the Nash-Sutcliffe efficiency (NSE; Nash and Sutcliffe 1970) and snowpack percent error (SPE), $\mathrm{MAE}/\bar{z}+$, where $\bar{z}+$ is the mean nonzero height.
@@ 3.1 Depth Timeseries
Figure 4: Performance of $M$ and $\tilde{M}$ against Snow17 with (SN17O) and without (SN17) observational $\mathrm{SWE}$ data for generating $z$ timeseries over the 44 validation sites. RMSE indicates root mean square error, NSE the Nash-Sutcliffe Efficiency, and SPE the average L1 error normalized to the average nonzero depth, to measure percent error. Boxes outline the 25% to 75% quantiles, with the bar at the median, while whiskers mark the extremes and dots indicate outliers, which lie beyond 1.5 times the interquartile range (box width) from the box.
@@ 3.4 Physical Behavior of Model
For SNOTEL instances with $T_{\mathrm{air},i}<0$, $\tilde{M}_{\mathrm{SWE}}$ violated the $d\mathrm{SWE}/dt\leq P_{\mathrm{snow}}$ condition only 0.6% of the time, with all violations under 1 mm/day and 95% under 0.75 mm/day. About half the instances showed $d\mathrm{SWE}/dt<P$ (prescribed bounds yielded equivalence in the rest), but the mean residual of these was -2 mm/day, with 95% above -5 mm/day, aligning with typical sublimation rates (Spehlmann et al., 2023; Liu et al., 2024) and effects not explicitly modeled. Given the SNOTEL data precision (2.54 mm) and negative average temperatures obscuring instantaneous positive temperatures, these results suggest strong adherence to mass conservation, even without explicitly enforcing it (as simple as choosing $f_{+}=P_{\mathrm{snow}}$ from section 22.2 for $\tilde{M}_{\mathrm{SWE}}$), highlighting the model’s ability to learn physical representations from data beyond those prescribed.
@@ 3.6 Finer Resolution Predictions
Time units only appear as rates in $M$ via precipitation, the lower bound, and the output. By predicting rates $dz/dt$ instead of accumulated $dz$, $M$ can be evaluated at varying time-steps without retraining. This flexibility allows testing at resolutions up to data limits, which for Kühtai is a maximum resolution of 15-minute intervals. While sub-hourly time resolutions are rarer in larger earth models, they are used in land models and will become increasingly frequent as research advances in fine-scale land-atmosphere interactions (Schär et al., 2020; Ban et al., 2021). Similarly, while multiday resolutions are rare, data availability can necessitate their use. Benchmarking across scales above and below the daily resolution used for training offers insights into the model’s limits and potential applications. However, beyond a week, average input variables do not effectively capture critical input dynamics (e.g., monthly average $T_{\mathrm{air}}$ would not reliably indicate time above freezing), leading to suppressed variations that undermine meaningful outputs, and alternative model choices would be more appropriate. Results spanning 15-minute to weekly resolutions are shown in Fig. 12 and Table 2.
@@
Figure 12: Output of $M$ at different resolutions for a subset of the site data from Kühtai, Austria. The graph overlays outputs at different resolutions for direct comparison.
@@ Acknowledgements.
Many code packages and languages cannot compute gradients for logical branches or min/max functions. This structure bypasses that limitation, allowing constraints to be enforced during training, even on legacy systems with minimal functionality. This mitigates the need to penalize the loss function, which can impede learning of the main objective (Rahimi et al., 2022). Enforcing constraints during training enables gradients and weight updates better-suited for predicting values within the boundaries. Unlike other penalty-free approaches like projecting outputs into constrained spaces, this method does not require constant or predefined thresholds and can explicitly predict boundary values rather than asymptotically approach them, which can inhibit learning by reducing gradient magnitudes (LeCun et al., 2012). Constraints under this construction can be analytically defined or parameterized and “learned” from data during training, even simultaneously with the predictive component, predicting both thresholds and values for entirely data-driven modeling. These functional forms can also be combined and stacked in larger networks or applied to non-network models. Figure 13 depicts the threshold constraint layers enveloping the predictive component from section 22.2, but they could be placed inside larger networks, layered, or stacked as part of a larger predictive model, or applied on any non-network model.
@@
Phenomena like wind drift can confound snow pillow data, as strong winds away from the wind sensor can push snow onto the pillow, creating positive measured $dz/dt$ with no precipitation and insufficient wind speeds for drift, as observed by Meyer et al. (2012) in some SNOTEL sites. Our constraints in section 22.2 can be violated by such events. However, only 1.5% of the training data and 2.2% of the testing data showed $dz/dt>0$ without precipitation, and established models like Crocus and SNOWPACK also do not account for positive growth due to wind redistribution when run standalone. These models instead focus on the compaction or erosion of snow by wind (Vionnet et al., 2012; Lehning et al., 2002); destructive effects our model captures. Ongoing work aims to integrate redistribution effects into snow models, and as this paper serves to introduce a framework compared to existing parameterizations, we find accounting for such effects beyond the scope of this study, representing an exciting area for future research. We invite interested parties to revise variables, thresholds, and datasets to the benefit of the community, as we envision broader adoption of this proposal beyond our specific formulation and calibration.

### v2

@@ 2.1 Overview
Our model choices leverage ML for seasonal snow simulation in climate models, prioritizing generalizability and computational efficiency. Contemporary paradigms in hydrology research focus on parameterizing SWE from $z$ and other data to constrain its value over global grids. However, within global climate models, SWE evolution is already well-constrained by explicitly implemented physical laws enforcing mass conservation and relatively well-understood fluxes such as sublimation, precipitation, and melt. The subsequent conversion of simulated SWE to variables like $z$ or $\rho_{\mathrm{snow}}$ is typically left to longstanding parameterizations. These parameterizations can be pre-calibrated offline prior to use in a snow model, or are often further tuned “online” within a snow model. The data used for calibration can be indirect (non-SWE data), such as energy and water fluxes or land surface albedo, or be direct measurements of snow depth or density, or even gridded SWE estimates. The resulting global simulations are sometimes employed in refining gridded SWE for calibrating other models, creating circular estimation and biases.
@@ 2.2 Model Structure
Figure 1: Structure of the model’s predictive component. Blue lines indicate a trainable linear transformation of the input (of $k$ scalar variables), including a bias. Colors indicate the activation function used upon collection at the node (ReLU = Rectified Linear Unit, ELU = Exponential Linear Unit, Id = Identity), as noted in the legend. The hyperparameter $n$ sets the width of the internal mixing layer.
@@ 2.3 Threshold Constraints for Snowpack Prediction
Constraints for $dz/dt$ should keep the depth tendency within physical limits to enable generalizability and stability when $M$ is integrated over time. This initial study selected the following basic constraints:
@@
These constraints can be expressed as threshold functions, the lower as $f_{-}=-z/\Delta t$ and the upper as $f_{+}={\rm{ReLU}}(p)\times\mathbf{1}_{\{P_{\mathrm{snow}}>0\}}$, where $p$ is the output of the predictive component and $\mathbf{1}$ is the indicator function. For these choices, $f_{-}$ is nonpositive and $f_{+}$ is nonnegative, which simplifies the constraint layer structure (see Appendix A), resulting in a final structure for $M$ as depicted in Fig. 2.
@@
The first constraint includes the time step $\Delta t$, but this does not explicitly affect the time dependency or resolution of the parameterization. The predictive component contains no time nor time-step dependence. Adjusting $\Delta t$ scales the constraint appropriately without altering the predictive component’s output, enabling the model’s use in adaptive time-step schemes (re-scaling the constraint per time step). This means the model in principle only requires training with data at one temporal resolution, though we would anticipate improved time-step independence if training data incorporated varied time intervals, or $dz/dt$ values in the minimum range anticipated during usage.
@@ 2.4 Data
We required training data with simultaneous snow and meteorological measurements, preferring collocated primary ground observations over reanalysis data due to known discrepancies (Meyer et al., 2023). One primary source is the US Snow Telemetry Network (SNOTEL) by the Natural Resources Conservation Service (NRCS). Data from 44 total SNOTEL sites in the contiguous United States (CONUS) have simultaneous availability of the necessary inputs, of which the entire reporting history until February 1st, 2024 was collected. These sites span diverse climates (see Fig 3), enhancing generalizability. For testing, 7 total Alaskan SNOTEL sites had the necessary inputs and were similarly collected, plus standard evaluation data from Kühtai, Austria (Krajči et al., 2017), Col de Porte, France (Lejeune et al., 2019), and the Reynolds Mountain East catchment (Reba et al., 2011) to assess the $M$’s ability to generalize to out-of-sample data. We additionally used data from Sodankylä, Finland (Essery et al., 2016), the upper Rofental (Warscher et al., 2024), and Yala Basecamp (Stigter et al., 2021; Shea et al., 2015) in the Himalayas to test performance across uncalibrated elevations and climate types.
@@ 2.6 Testing
The integer $K$ specifies sequential data transitions ($K=1$) or “gaps” to traverse in timeseries data ($K>1$) due to missing or cleaned data. $M$’s built-in constraints ensure non-negative $z$ values when the step size is the designated $\Delta t$ (i.e. when $K=1$; for providing other $\Delta t$ see Section 3f), though this choice of time-stepping procedure can create negative $z$ when $K>1$ (rates constrained for step-size $\Delta t$ are applied over $K\Delta t>\Delta t$). In such cases, negative $z$ values were set to zero, and similarly, timeseries were “reset” to observed values $\hat{z}_{i+K}=z_{i+K}$ whenever $K>K_{\mathrm{max}}=5$ days, to avoid attributing error from the method choice to $M$ for fair evaluation. When $M$ is used within a global model simulation, such gaps would not occur since the inputs would be available at every step, allowing all $z_{i}$ to obey prescribed bounds. Selecting $K_{\mathrm{max}}>1$ day additionally reduced the number of resets that would otherwise beneficially skew performance metrics, with $K_{\mathrm{max}}=5$ ensuring resets in less than 2.3% of cases (median frequency 0.19%), with many over whole years or no-snow periods (not impacting snow simulation). Evaluation metrics included root mean square error (RMSE), mean absolute error (MAE), bias (B), and median percent error (MPE, on $dz/dt$ for regression and $z$ for timeseries). For timeseries, we included the Nash-Sutcliffe efficiency (NSE; Nash and Sutcliffe 1970) and snowpack percent error (SPE), $\mathrm{MAE}/\bar{z}+$, where $\bar{z}+$ is the mean nonzero height.
@@ 3.1 Depth Timeseries
Figure 4: Performance of $M$ and $\tilde{M}$ against Snow17 with (SN17O) and without (SN17) observational $\mathrm{SWE}$ data for generating $z$ timeseries over the 44 validation sites. RMSE indicates root mean square error, NSE the Nash-Sutcliffe Efficiency, and SPE the average L1 error normalized to the average nonzero depth, to measure percent error. Boxes outline the 25% to 75% quantiles, with the bar at the median, while whiskers mark the extremes and dots indicate outliers, which lie beyond 1.5 times the interquartile range (box width) from the box. Vertical axes limits are chosen to show all data points.
@@ 3.4 Physical Behavior of Model
For SNOTEL instances with $T_{\mathrm{air},i}<0$, $\tilde{M}_{\mathrm{SWE}}$ violated the $d\mathrm{SWE}/dt\leq P_{\mathrm{snow}}$ condition only 0.6% of the time, with all violations under 1 mm/day and 95% under 0.75 mm/day. About half the instances showed $d\mathrm{SWE}/dt<P_{\mathrm{snow}}$ (prescribed bounds yielded equivalence in the rest), but the mean residual of these was -2 mm/day, with 95% above -5 mm/day, aligning with typical sublimation rates (Spehlmann et al., 2023; Liu et al., 2024) and effects not explicitly modeled. Given the SNOTEL data precision (2.54 mm) and negative average temperatures obscuring instantaneous positive temperatures, these results suggest strong adherence to mass conservation, even without explicitly enforcing it (as simple as choosing $f_{+}=P_{\mathrm{snow}}$ from Section 22.2 for $\tilde{M}_{\mathrm{SWE}}$), highlighting the model’s ability to learn physical representations from data beyond those prescribed.
@@ 3.6 Finer Resolution Predictions
Time units only appear as rates in $M$ via precipitation, the lower bound, and the output. By predicting rates $dz/dt$ instead of accumulated $dz$, $M$ can be evaluated at varying time-steps without retraining, by merely resetting the constraint function’s scaling of $1/\Delta t$ without altering the trained predictive weights (which were trained only once in section 3a, with $\Delta t$ = 1 day). This flexibility allows testing at resolutions up to data limits, which for Kühtai is a maximum resolution of 15-minute intervals. While sub-hourly time resolutions are rarer in larger earth models, they are used in land models and will become increasingly frequent as research advances in fine-scale land-atmosphere interactions (Schär et al., 2020; Ban et al., 2021). Similarly, while multiday resolutions are rare, data availability can necessitate their use. Benchmarking across scales above and below the daily resolution used for training offers insights into the model’s limits and potential applications. However, beyond a week, average input variables do not effectively capture critical input dynamics (e.g., monthly average $T_{\mathrm{air}}$ would not reliably indicate time above freezing), leading to suppressed variations that undermine meaningful outputs, and alternative model choices would be more appropriate. Results of repeating the timeseries generation with $M$ as described in Section 3f for the Kühtai data with different values of $\Delta t$ in the constraint but identical weights otherwise, spanning 15-minute to weekly resolutions, are shown in Fig. 12 and Table 2.
@@
Figure 12: Output of $M$ at different resolutions for a subset of the site data from Kühtai, Austria. The graph overlays outputs at different resolutions for direct comparison. Sub-hourly (15 min, 30 min) curves are nearly indistinguishable from the hourly curve.
@@ Acknowledgements.
Many code packages and languages cannot compute gradients for logical branches or min/max functions. This structure bypasses that limitation, allowing constraints to be enforced during training, even on legacy systems with minimal functionality. This mitigates the need to penalize the loss function, which can impede learning of the main objective (Rahimi et al., 2022). Enforcing constraints during training enables gradients and weight updates better-suited for predicting values within the boundaries. Unlike other penalty-free approaches like projecting outputs into constrained spaces, this method does not require constant or predefined thresholds and can explicitly predict boundary values rather than asymptotically approach them, which can inhibit learning by reducing gradient magnitudes (LeCun et al., 2012). Constraints under this construction can be analytically defined, or even parameterized and “learned” from data during training, even simultaneously with the predictive component (e.g., a network for prediction and a network for boundary values, trained simultaneously), predicting both thresholds and values for entirely data-driven modeling. These functional forms can also be combined and stacked in larger networks or applied to non-network models. Figure 13 depicts the threshold constraint layers enveloping the predictive component from section 22.2, but they could be placed inside larger networks, layered, or stacked as part of a larger predictive model, or applied on any non-network model.
@@
Phenomena like wind drift can confound snow pillow data, as strong winds away from the wind sensor can push snow onto the pillow, creating positive measured $dz/dt$ with no precipitation and insufficient wind speeds for drift, as observed by Meyer et al. (2012) in some SNOTEL sites. Our constraints in section 22.2 can be violated by such events. However, only 1.5% of the training data and 2.2% of the testing data showed $dz/dt>0$ without precipitation, and established models like Crocus and SNOWPACK also do not account for positive growth due to wind redistribution when run standalone. These models instead focus on the compaction or erosion of snow by wind (Vionnet et al., 2012; Lehning et al., 2002); destructive effects our model captures. Ongoing work aims to integrate redistribution effects into snow models, and as this paper serves to introduce a framework compared to existing parameterizations, we find accounting for such effects beyond the scope of this study, representing an exciting area for future research. We invite interested parties to revise the variables, thresholds, and datasets in this initial formulation to the benefit of the community, as we envision broader adoption of this proposal beyond our specific formulation and calibration.